Suppose that capital costs $50 per unit and labour costs $20 per unit. If the marginal product of capital is 100 and the marginal product of labour is 30, a cost-minimizing firm should

employ more of both capital and labour.

employ less of both capital and labour.

employ more capital and less labour.

employ less capital and more labour.

not change its current factor use.
Answer #1

Answer

cost is minimum at the production level where the marginal product per dollar is equal.

MP per dollar =MP/input price

MP per dollar of labor =30/20=0.5

MP per dollar of capital =100/50=2

the firm should increase capital input and decrease labor input up to MP per dollar is equal

Option 3

employ more capital and less labour